# Higher Order Dualities between Prime Ideals

### v1

@@
Abstract: Extending the works of Alladi [Al77] and Sweeting and Woo [SW18], we state and prove the general higher order duality between prime ideals in number rings. We then use the second order duality to obtain the a new formula for the Chebotarev Density involving sums of the generalized Möbius function and the prime ideal counting function. We also provide two estimates of such sums as an application of the duality identity. A discussion of the duality in a slightly more general setting is done at the end.
Keywords: Duality between prime ideals, generalized Möbius function, number of prime ideals, largest and smallest norm of prime ideals, Chebotarev Density, Chebotarev Density Theorem, Galois extensions, Prime Ideal Theorem, Artin Symbol
@@ 1 Introduction and Background
Since time immemorial, the Möbius function have been an integral and essential part of analytic number theory and several important results and theories have come up around it. Let us first recall this significantly important function defined on positive integers:
@@
One of the most well-known examples of the use of the Möbius function is in its relation with the Riemann Zeta function. Using the Euler product formula of $\zeta(s)$, one can show that the Dirichlet series with $\mu(n)$ as its coefficients is the reciprocal of the zeta function. Further,
$\sum_{n=1}^{\infty}\frac{\mu(n)}{n}=\lim_{s\to 1^{+}}\frac{1}{\zeta(s)}=0$
@@
is (elementarily) equivalent to the Prime Number Theorem. A surprising generalization of the (1.1) was observed by Alladi [Al77], in 1977, as a consequence of the following Duality Lemma he proved in the same paper. He proved that if $f$ is a function defined on primes, then:
@@
where $P_{1}(n)$ and $p_{1}(n)$ are the largest and the smallest prime factors of the integer $n$ respectively. Using (1.2) and some properties of the Möbius function, Alladi showed that if $f$ is a bounded function on the primes such that
@@
and vice-versa. We notice that (1.1) is indeed a special case of (1.4). If we choose $f$ to be an arithmetic function on primes defined as $f(p)=1$ for all primes $p$, then such an $f$ will yield from (1.3) that
@@
for all positive integers $k,\ell$, such that $1\leq\ell<k$ and $(\ell,k)=1$. We note here that the sum on the left of (1.6) gives us a way of slicing the sum on the left of (1.5) into $\varphi(k)$ subseries, each of them converging to the same value! Alladi gave a quantitative proof of (1.6) in [Al77] which involves the use of the duality identity (1.2).
@@
$\sum_{\begin{subarray}{c}n=2\\ \left[\frac{K/\mathbb{Q}}{p_{1}(n)}\right]\end{subarray}}^{\infty}\frac{\mu(n)}{n}=-\frac{|C|}{|G|}$ (1.7)
@@
Extending the previous works by Alladi and Dawsey, in 2018, Sweeting and Woo [SW18] proved similar density type results by making the choice of the base field of the Galois extension to be any arbitrary number field. In particular, if $C\subset G=Gal(L/K)$ is a conjugacy class, then they proved the following new formula for the Chebotarev Density $\frac{|C|}{|G|}$:
@@
where $\mu_{K}(I)$ denotes the generalized Möbius function (see below) and $S(L/K;C)$ is the set of ideals $I\in\mathcal{O}_{K}$ such that $I$ has a unique prime divisor $\mathfrak{p}_{1}$ of minimum norm and the Artin sumbol $\left[\frac{L/K}{\mathfrak{p}}\right]$ is $C$. We would like to note here that proving (1.8) required new machinery as Alladi’s Duality lemma is only valid for the classical case. Since Dawsey proves her results by considering the base field to be $\mathbb{Q}$, Alladi Duality works fine as it is applicable to integers, which act like the ideals in $\mathcal{O}_{\mathbb{Q}}$, i.e. $\mathbb{Z}$. Hence, Sweeting and Woo prove the following first order duality between prime ideals (see Lemma 2.1 in [SW18]): if $f(I)$ denote the indicator function of $S(L/K;C)$, then for ideals $I$, we have
$\sum_{J\supset I}\mu(J)f(J)=-Q_{C}(J),$ (1.9)
where, the definition of the quantity in the RHS above is given by
@@
where $M(I)=\max_{I\subset\mathfrak{p}}N(\mathfrak{p})$. This duality between the prime ideals with smallest and largest norms proved crucial in proving (1.8). Not only that, the proof of (1.8) is also of quantitative in nature, which makes it even more interesting. Several other results in [SW18] are extensions of those in both [Al77] and [Da17], hence connecting the ideas of both the classical case and the algebraic case. The most important one to take a note here is Theorem 1.1, where they use the strong form of CDT. One can indeed find a correspondence between the same and the uniform distribution of the largest prime factors in an arithmetic progression $\ell\;(mod\;k)$ [Al77] and the existence of the density of $f(P_{1}(n))$, where $f$ is the characteristic function of the prime numbers satisfying $\left[\frac{K/\mathbb{Q}}{\mathfrak{p}}\right]=C$ [Da17].
@@
$\sum_{2\leq n\leq x}f(P_{1}(n))\sim\kappa x\quad\text{and}\quad\sum_{2\leq n\leq x}f(P_{2}(n))\sim\kappa x$ (1.11)
@@
We note here that $\omega(n)$ is defined by the number of distinct prime factors of $n$ and $P_{2}(n)$ is defined to be the second largest prime factor on $n$. In both [AJ24] and [Se25], there have been discussions on the definitions of the second largest prime factors. Following them, $P_{2}(n)$ here is defined by the largest prime factor of $n$ strictly less than $P_{1}(n)$. Their paper itself gives us the two examples that help us understand the above Theorem 13. More enchantingly, both of those examples have been proved as theorems (see Theorem 4(i) and Theorem 10 in [AJ24]) and the proofs are quantitative in nature.
@@
$\sum_{\begin{subarray}{c}n=2\\ \left[\frac{K/\mathbb{Q}}{p}\right]=C\end{subarray}}^{\infty}\frac{\mu(n)\omega(n)(\omega(n)-1)}{n}=\frac{|C|}{|G|},$
providing us a new formula of the Chebotarev Density.
Like Dawsey, [Se25] also extends results in [AJ24] to the setting of finite Galois extensions with the base field chosen to be the field of rationals $\mathbb{Q}$. With the work of Sweeting and Woo in the picture as discussed above, it is quite natural to study the generalization of results of Alladi and Johnson, and the author himself, to the setting of arbitrary finite Galois extensions. In that case too, one would need a new version of duality between prime ideals that involves prime ideals of $k^{th}$ largest norms. In this paper, we first state and prove (see §2, Theorem 2.4) the higher order duality lemma between prime ideals. We then use the second order duality (see Theorem 2.3) specifically to prove the density type extensions222new formula for the Chebotarev Density of results in [AJ24] and [Se25]. Sections that will follow the proofs of our main theorems will also provide applications of this new general duality between prime ideals in number fields and remark on how it can prove to be essential in generalizing all such duality related results that have been derived by several other researchers (see §5,§6) in the classical case. For now, let us look at the necessary notations and preliminary theorems that we will use in the proofs to come.
Throughout rest of the paper, $L/K$ will denote a finite Galois extension of number fields such that $G=Gal(L/K)$. The corresponding rings of integers of the fields involved in this extension will be denoted as $\mathcal{O}_{K}$ and $\mathcal{O}_{L}$. Let $\mathfrak{p}\in\mathcal{O}_{K}$ denote a prime ideal in $\mathcal{O}_{K}$. Then the ideal generated by $\mathfrak{p}$ in $\mathcal{O}_{L}$ can be decomposed into a product of distinct prime ideals $\mathfrak{p}_{i}$’s in $\mathcal{O}_{L}$ lying above $\mathfrak{p}$, i.e.
@@
where $e_{i}$’s are positive integers. Our only concern, in the results that follow, are primes ideals in $\mathcal{O}_{K}$ that are unramified in $\mathcal{O}_{L}$, i.e. the corresponding integers $e_{i}$, for each $i$, are 1. We further recall that the absolute norm of a non-zero ideal $\mathcal{I}$ in any number ring $\mathcal{O}$ is given by
@@
$\mathfrak{P}_{C}:=\left\{\mathfrak{p}\in\mathfrak{P}(K):\;\mathfrak{p}\text{ is unramified in $L$ and }\left[\frac{L/K}{\mathfrak{p}}\right]=C\right\}$
@@
$\pi_{C}(X,L/K)=\frac{|C|}{|G|}\cdot\frac{X}{\log X}+o\left(\frac{X}{\log X}\right)$
@@
$\left|\pi_{C}(X,L/K)-\frac{|C|}{|G|}Li(X)\right|\ll Xe^{-c_{1}\sqrt{\frac{\log x}{n_{L}}}}$
@@
It is important to note the error estimate above in Theorem 1.2 as it is a crucial element in the proofs to come. We also note Landau’s form of the Prime Ideal Theorem [5 in SW] which will also be essential later on. We denote $\pi(K;X)$ to be the number of prime ideals in $\mathcal{O}_{K}$ whose norm is bounded by $X$.
@@
Theorem 1.3 is the more explicit version of the Prime Ideal Theorem by Landau. It is also worthwhile to note the asymptotic version of the same (see also [MV07], p. 194), i.e. if $K$ is a number field then
$\pi(K;X)\sim\frac{X}{\log X}$
@@
In the next section, we will state and prove the general higher order duality between prime ideals. We will first look at the second order duality to understand how the proof works and then move ahead with the general proof. Note that Sweeting and Woo proved the first order duality (see Lemma 2.1 in [SW18]) and we will see how it is consistent with our general duality identities. In the sections that follow §2, we use the second order duality to get a new formula for the Chebotarev Density, with their quantitative proofs.
@@ 2 General Duality Between Prime Ideals
$\sum_{d|n}\mu(d)f(P_{k}(d))=(-1)^{k}{\omega(n)-1\choose k-1}f(p_{1}(n))$
$\sum_{d|n}\mu(d)f(p_{k}(d))=(-1)^{k}{\omega(n)-1\choose k-1}f(P_{1}(n))$
@@
$\sum_{d|n}\mu(d){\omega(d)-1\choose k-1}f(P_{1}(d))=(-1)^{k}f(p_{k}(n))$
$\sum_{d|n}\mu(d){\omega(d)-1\choose k-1}f(p_{1}(d))=(-1)^{k}f(P_{k}(n))$
As one can observe, this is applicable only for integers and therefore, we need and will proceed to state and prove the duality version for prime ideals in arbitrary number fields. Let us first define a few quantities and sets essential for stating the duality identities. Let $N(I)$ denote the norm of an ideal $I$. For an ideal $I\subset\mathcal{O}_{K}$, we define $M_{k}(I)$ to be the norm of the prime ideals in the decomposition of the ideal $I$ with the $k^{th}$ largest norms. Of course, the quantity $M_{k}(I)$ makes sense only if $I$ is contained in prime ideals with $k$ many distinct norms. Note that there might not be a unique prime ideal containing $I$ with a particular norm. In view of this fact, we define the sets $Q^{k}(I)$ and $Q_{C}^{k}(I)$ as follow: for a finite Galois extension $L/K$ and given a fixed conjugacy class $C\subset G=Gal(L/K)$,
$Q_{C}^{k}(I)$ $:=\#\left\{\mathfrak{p}:\;I\subset\mathfrak{p};\;M_{k}(I)=N(\mathfrak{p})\;\text{and}\;\left[\frac{L/K}{\mathfrak{p}}\right]=C,\;\mathfrak{p}\;\text{is unramified}\right\}$
$Q^{k}(I)$ $:=\#\{\mathfrak{p}:\;I\subset\mathfrak{p};\;M_{k}(I)=Nm(\mathfrak{p})\}\hskip 85.35826pt$
From now onwards, we only focus on such ideals $I$ that have a unique prime ideal of the smallest norm. For a fixed $I$ satisfying such a property, we denote the unique prime factor of $I$ with the smallest norm as $\mathfrak{p}_{1}$. Following the convention proposed in [SW18], we also call these ideals to be the salient ideals in $\mathcal{O}_{K}$. Let us define the set $S(L/K;C)$ as follows:
$S(L/K;C):=\left\{I\subset\mathcal{O}_{K}:\;I\text{ is salient and }\left[\frac{L/K}{\mathfrak{p}_{1}}\right]=C,\;\text{where $\mathfrak{p}_{1}$ is unramified}\right\}$
@@
Note that as per our definitions, $Q_{C}(I)$ and $Q_{C}^{1}(I)$ denote the same set. We further use $\omega_{K}(n)$ to denote the prime ideal counting function
$\omega_{K}(n)=\sum_{\mathfrak{p}\supset I}1$
We first state and prove what we view as the second order duality between prime ideals, i.e. duality for $k=2$.
@@
Let $K$ be a number field and $I\subset\mathcal{O}_{K}$ be an ideal. Let $f$ be the characteristic function of the set $S(L/K;C)$. Then for ideals satisfying $Q^{1}(I)=1$555this means that such ideals have a unique prime ideal factor with the maximum norm, the following identity holds:
@@
We start with a fixed ideal $I$ satisfying the required hypothesis, such that $I=\mathfrak{p}_{1}^{e_{1}}\mathfrak{p}_{2}^{e_{2}}...\mathfrak{p}_{r}^{e_{r}}$. Without the loss of generality, we assume that in the factorization above, $N(\mathfrak{p}_{i})\leq N(\mathfrak{p}_{i+1})$, that is, we factor $I$ into prime ideals in an ascending order of their norms. We rewrite the factorization as $I=I_{1}I_{2}...I_{s}$, where each of these $I_{j}$’s are product of the respective prime ideals of the same norm. For example, if $\mathfrak{p}_{i_{1}}^{e_{i_{1}}}\mathfrak{p}_{i_{2}}^{e_{i_{2}}}\supset I_{i}$, for some integers $i,i_{1},i_{2}$, then $N(\mathfrak{p}_{i_{1}})=N(\mathfrak{p}_{i_{2}})$. Therefore,
@@
$=$ $\sum_{j=1}^{s-1}\sum_{\begin{subarray}{c}J\supset I_{j}\\ J\not\supset I_{k^{\prime}},\;\forall k^{\prime}\geq j+1\\ J\neq\mathcal{O}_{K}\end{subarray}}\mu_{K}(J)f(J)\sum_{J^{\prime}\supset I_{j+1}...I_{s}}\mu_{K}(J^{\prime})(\omega_{K}(J)-1+\omega_{K}(J^{\prime}))$ (2.1)
We note that in the above equality, the outer sum on the right ranges to s1s-1, since the contribution from the ideal $I_{s}$ is 0. Also, we do not need the term $f(J^{\prime})$ as by the definition of $f$, we only need to focus on the ideal with the smallest norm, which of course contains $J$ and not $J^{\prime}$. Therefore, continuing from (2.1), we have
@@
$+\sum_{j=1}^{s-1}\sum_{\begin{subarray}{c}J\supset I_{j}\\ J\not\supset I_{k^{\prime}},\;\forall k^{\prime}\geq j+1\\ J\neq\mathcal{O}_{K}\end{subarray}}\mu_{K}(J)f(J)\sum_{J^{\prime}\supset I_{j+1}...I_{s}}\mu_{K}(J^{\prime})\omega_{K}(J^{\prime})=:\mathcal{S}_{1}+\mathcal{S}_{2}$ (2.2)
@@
$(1-x)^{t}=\sum_{i=0}^{t}(-1)^{i}{t\choose i}x^{i}$ (2.4)
Hence, the RHS of (2.4) with $x=1$ yields the RHS of (2.3), and therefore, we have from both the equations that
$\sum_{J^{\prime}\supset I_{j+1}...I_{s}}\mu_{K}(J^{\prime})=\begin{cases}1,\;\text{when $I_{j+1}...I_{s}=\mathcal{O}_{k}$}\\ 0,\;\text{otherwise}\end{cases}$ (2.5)
From (2.5), it is clear that the only contribution from $\mathcal{S}_{1}$ is when $I_{j+1}...I_{s}=\mathcal{O}_{K}$, which is again only possible if $j=s$. But that is not possible since $j$ ranges from $1$ to s1s-1 and hence, from (2.5), we conclude that
$\mathcal{S}_{1}=0$ (2.6)
@@
$\sum_{J^{\prime}\supset I_{j+1}...I_{s}}\mu_{K}(J^{\prime})\omega_{K}(J^{\prime})=\sum_{i=1}^{t}(-1)^{i}{t\choose i}i$ (2.7)
@@
$\mathcal{S}_{2}=-\sum_{\begin{subarray}{c}J\supset I_{s-1}\\ J\not\supset I_{s}\\ J\neq\mathcal{O}_{K}\end{subarray}}\mu_{K}(J)f(J)=Q_{C}^{2}(I)$ (2.10)
To understand the last equality, we observe that $I_{s-1}$ is a product of prime ideals containing $I$ with the second largest norm. Now, the only positive contribution from the terms in the sum in (2.10) are when we choose $J$ to just be the prime ideals, since:

### v2

@@
Abstract: Extending the works of Alladi [Al77] and Sweeting and Woo [SW18], we state and prove the general higher order duality between prime ideals in number rings. We then use the second order duality to obtain a new formula for the Chebotarev Density involving sums of the generalized Möbius function and the prime ideal counting function. We also provide two estimates of such sums as an application of the duality identity. A discussion of the duality in a slightly more general setting is done at the end.
Keywords: Duality between prime ideals, generalized Mobius function, number of prime ideals, largest and smallest norm of prime ideals, Chebotarev Density, Chebotarev Density Theorem, Galois extensions, Prime Ideal Theorem, Artin Symbol
@@ 1 Introduction and Background
Since the $19^{th}$ century, the Möbius function has been an integral and essential part of analytic number theory, and several important results and theories have come up around it. Let us first recall this significantly important function defined on positive integers:
@@
One of the most well-known examples of the use of the Möbius function is in its relation with the Riemann Zeta function. Using the Euler product formula of $\zeta(s)$, one can show that the Dirichlet series, with $\mu(n)$ as its coefficients, is the reciprocal of the zeta function. Further,
$\sum_{n=1}^{\infty}\frac{\mu(n)}{n}=\lim_{s\to 1^{+}}\frac{1}{\zeta(s)}=0.$
@@
is (elementarily) equivalent to the Prime Number Theorem. A surprising generalization of (1.1) was observed by Alladi [Al77] in 1977, as a consequence of the following Duality Lemma he proved in the same paper. He proved that if $f$ is a function defined on primes, then:
@@
where $P_{1}(n)$ and $p_{1}(n)$ are the largest and the smallest prime factors of the integer $n$, respectively. Using (1.2) and some properties of the Mobius function, Alladi showed that if $f$ is a bounded function on the primes such that
@@
and vice versa. We notice that (1.1) is indeed a special case of (1.4). If we choose $f$ to be an arithmetic function on primes defined as $f(p)=1$ for all primes $p$, then such an $f$ will yield from (1.3) that
@@
for all positive integers $k,\ell$, such that $1\leq\ell<k$ and $(\ell,k)=1$. We note here that the sum on the left of (1.6) gives us a way of slicing the sum on the left of (1.5) into $\varphi(k)$ subseries, each of them converging to the same value! Alladi gave a quantitative proof of (1.6) in [Al77], which involves the use of the duality identity (1.2).
@@
$\sum_{\begin{subarray}{c}n=2\\ \left[\frac{K/\mathbb{Q}}{p_{1}(n)}\right]=C\end{subarray}}^{\infty}\frac{\mu(n)}{n}=-\frac{|C|}{|G|}.$ (1.7)
@@
Extending the previous works by Alladi and Dawsey, in 2018, Sweeting and Woo [SW18] proved similar density-type results by making the choice of the base field of the Galois extension to be any arbitrary number field. In particular, if $C\subset G=Gal(L/K)$ is a conjugacy class, then they proved the following new formula for the Chebotarev Density $\frac{|C|}{|G|}$:
@@
where $\mu_{K}(I)$ denotes the generalized Möbius function (see below) and $S(L/K;C)$ is the set of ideals $I\in\mathcal{O}_{K}$ such that $I$ has a unique prime divisor $\mathfrak{p}_{1}$ of minimum norm and the Artin sumbol $\left[\frac{L/K}{\mathfrak{p}}\right]$ is $C$. We would like to note here that proving (1.8) required new machinery, as Alladi’s Duality lemma is only valid for the classical case. Since Dawsey proves her results by considering the base field to be $\mathbb{Q}$, Alladi Duality works fine as it is applicable to integers, which act like the ideals in $\mathcal{O}_{\mathbb{Q}}$, i.e., $\mathbb{Z}$. Hence, Sweeting and Woo prove the following first order duality between prime ideals (see Lemma 2.1 in [SW18]): if $f(I)$ denotes the indicator function of $S(L/K;C)$, then for ideals $I$, we have
$\sum_{J\supset I}\mu_{K}(J)f(J)=-Q_{C}(J),$ (1.9)
where the definition of the quantity in the RHS above is given by
@@
where $M(I)=\max_{I\subset\mathfrak{p}}N(\mathfrak{p})$. This duality between the prime ideals with the smallest and largest norms proved crucial in proving (1.8). Not only that, the proof of (1.8) is also quantitative in nature, which makes it even more interesting. Several other results in [SW18] are extensions of those in both [Al77] and [Da17], hence connecting the ideas of both the classical case and the algebraic case. The most important one to take note of here is Theorem 1.1, where they use the strong form of CDT. One can indeed find a correspondence between the same and the uniform distribution of the largest prime factors in an arithmetic progression $\ell\;(mod\;k)$ [Al77] and the existence of the density of $f(P_{1}(n))$, where $f$ is the characteristic function of the prime numbers satisfying $\left[\frac{K/\mathbb{Q}}{\mathfrak{p}}\right]=C$ [Da17].
@@
$\sum_{2\leq n\leq x}f(P_{1}(n))\sim\kappa x\quad\text{and}\quad\sum_{2\leq n\leq x}f(P_{2}(n))\sim\kappa x,$ (1.11)
@@
We note here that $\omega(n)$ is defined by the number of distinct prime factors of $n$ and $P_{2}(n)$ is defined to be the second largest prime factor of $n$. In both [AJ24] and [Se25], there have been discussions on the definitions of the second largest prime factors. Following them, $P_{2}(n)$ here is defined by the largest prime factor of $n$ strictly less than $P_{1}(n)$. Their paper itself gives us the two examples that help us understand the above Theorem 13. More enchantingly, both of those examples have been proved as theorems (see Theorem 4(i) and Theorem 10 in [AJ24]), and the proofs are quantitative in nature.
@@
$\sum_{\begin{subarray}{c}n=2\\ \left[\frac{K/\mathbb{Q}}{p}\right]=C\end{subarray}}^{\infty}\frac{\mu(n)(\omega(n)-1)}{n}=\frac{|C|}{|G|},$
providing us a new formula for the Chebotarev Density.
Like Dawsey, [Se25] also extends results in [AJ24] to the setting of finite Galois extensions with the base field chosen to be the field of rationals $\mathbb{Q}$. With the work of Sweeting and Woo in the picture as discussed above, it is quite natural to study the generalization of the results of Alladi and Johnson, and the author himself, to the setting of arbitrary finite Galois extensions. In that case, too, one would need a new version of duality between prime ideals that involves prime ideals of $k^{th}$ largest norms. In this paper, we first state and prove (see §2, Theorem 2.4) the higher order duality lemma between prime ideals. We then use the second order duality (see Theorem 2.3) specifically to prove the density type extensions222new formula for the Chebotarev Density of results in [AJ24] and [Se25]. Sections that will follow the proofs of our main theorems will also provide applications of this new general duality between prime ideals in number fields and remark on how it can prove to be essential in generalizing all such duality-related results that have been derived by several other researchers (see §5,§6) in the classical case. For now, let us look at the necessary notations and preliminary theorems that we will use in the proofs to come.
Throughout the rest of the paper, $L/K$ will denote a finite Galois extension of number fields such that $G=Gal(L/K)$. The corresponding rings of integers of the fields involved in this extension will be denoted as $\mathcal{O}_{K}$ and $\mathcal{O}_{L}$. Let $\mathfrak{p}\in\mathcal{O}_{K}$ denote a prime ideal in $\mathcal{O}_{K}$. Then the ideal generated by $\mathfrak{p}$ in $\mathcal{O}_{L}$ can be decomposed into a product of distinct prime ideals $\mathfrak{p}_{i}$’s in $\mathcal{O}_{L}$ lying above $\mathfrak{p}$, i.e.
@@
where $e_{i}$’s are positive integers. Our only concern, in the results that follow, is prime ideals in $\mathcal{O}_{K}$ that are unramified in $\mathcal{O}_{L}$, i.e., the corresponding integers $e_{i}$, for each $i$, are 1. We further recall that the absolute norm of a non-zero ideal $\mathcal{I}$ in any number ring $\mathcal{O}$ is given by
@@
$\mathfrak{P}_{C}:=\left\{\mathfrak{p}\in\mathfrak{P}(K):\;\mathfrak{p}\text{ is unramified in $L$ and }\left[\frac{L/K}{\mathfrak{p}}\right]=C\right\}.$
@@
$\pi_{C}(X,L/K)=\frac{|C|}{|G|}\cdot\frac{X}{\log X}+o\left(\frac{X}{\log X}\right).$
@@
$\left|\pi_{C}(X,L/K)-\frac{|C|}{|G|}Li(X)\right|\ll Xe^{-c_{1}\sqrt{\frac{\log x}{n_{L}}}},$
@@
It is important to note the error estimate above in Theorem 1.2 as it is a crucial element in the proofs to come. We also note Landau’s form of the Prime Ideal Theorem [La03], which will also be essential later on. We denote $\pi(K;X)$ to be the number of prime ideals in $\mathcal{O}_{K}$ whose norm is bounded by $X$.
@@
Theorem 1.3 is the more explicit version of the Prime Ideal Theorem by Landau. It is also worthwhile to note the asymptotic version of the same (see also [MV07], p. 194), i.e., if $K$ is a number field then
$\pi(K;X)\sim\frac{X}{\log X}.$
@@
In the next section, we will state and prove the general higher order duality between prime ideals. We will first look at the second order duality to understand how the proof works and then move ahead with the general proof. Note that Sweeting and Woo proved the first order duality (see Lemma 2.1 in [SW18]), and we will see how it is consistent with our general duality identities. In the sections that follow §2, we use the second order duality to get a new formula for the Chebotarev Density, with its quantitative proof.
@@ 2 General Duality Between Prime Ideals
$\sum_{d|n}\mu(d)f(P_{k}(d))=(-1)^{k}{\omega(n)-1\choose k-1}f(p_{1}(n)),$
$\sum_{d|n}\mu(d)f(p_{k}(d))=(-1)^{k}{\omega(n)-1\choose k-1}f(P_{1}(n)),$
@@
$\sum_{d|n}\mu(d){\omega(d)-1\choose k-1}f(P_{1}(d))=(-1)^{k}f(p_{k}(n)),$
$\sum_{d|n}\mu(d){\omega(d)-1\choose k-1}f(p_{1}(d))=(-1)^{k}f(P_{k}(n)).$
As one can observe, this is applicable only for integers, and therefore, we need and will proceed to state and prove the duality version for prime ideals in arbitrary number fields. Let us first define a few quantities and sets essential for stating the duality identities. Let $N(I)$ denote the norm of an ideal $I$. For an ideal $I\subset\mathcal{O}_{K}$, we define $M_{k}(I)$ to be the norm of the prime ideals in the decomposition of the ideal $I$ with the $k^{th}$ largest norms. Of course, the quantity $M_{k}(I)$ makes sense only if $I$ is contained in prime ideals with $k$ many distinct norms. Note that there might not be a unique prime ideal containing $I$ with a particular norm. In view of this fact, we define the sets $Q^{k}(I)$ and $Q_{C}^{k}(I)$ as follow: for a finite Galois extension $L/K$ and given a fixed conjugacy class $C\subset G=Gal(L/K)$,
$Q_{C}^{k}(I)$ $:=\#\left\{\mathfrak{p}:\;I\subset\mathfrak{p};\;M_{k}(I)=N(\mathfrak{p})\;\text{and}\;\left[\frac{L/K}{\mathfrak{p}}\right]=C,\;\mathfrak{p}\;\text{is unramified}\right\},$
$Q^{k}(I)$ $:=\#\{\mathfrak{p}:\;I\subset\mathfrak{p};\;M_{k}(I)=Nm(\mathfrak{p})\}.$
From now onwards, we only focus on such ideals $I$ that have a unique prime ideal of the smallest norm. For a fixed $I$ satisfying such a property, we denote the unique prime factor of $I$ with the smallest norm as $\mathfrak{p}_{1}$. Following the convention proposed in [SW18], we also call these ideals the salient ideals in $\mathcal{O}_{K}$. Let us define the set $S(L/K;C)$ as follows:
$S(L/K;C):=\left\{I\subset\mathcal{O}_{K}:\;I\text{ is salient and }\left[\frac{L/K}{\mathfrak{p}_{1}}\right]=C,\;\text{where $\mathfrak{p}_{1}$ is unramified}\right\}.$
@@
Note that as per our definitions, $Q_{C}(I)$ and $Q_{C}^{1}(I)$ denote the same set. We further use $\omega_{K}(I)$ to denote the prime ideal counting function
$\omega_{K}(I)=\sum_{\mathfrak{p}\supset I}1.$
We first state and prove what we view as the second order duality between prime ideals, i.e., duality for $k=2$.
@@
Let $K$ be a number field and $I\subset\mathcal{O}_{K}$ be an ideal. Let $f$ be the characteristic function of the set $S(L/K;C)$. Then, for ideals satisfying $Q^{1}(I)=1$555this means that such ideals have a unique prime ideal factor with the maximum norm, the following identity holds:
@@
We start with a fixed ideal $I$ satisfying the required hypothesis, such that $I=\mathfrak{p}_{1}^{e_{1}}\mathfrak{p}_{2}^{e_{2}}...\mathfrak{p}_{r}^{e_{r}}$. Without the loss of generality, we assume that in the factorization above, $N(\mathfrak{p}_{i})\leq N(\mathfrak{p}_{i+1})$, that is, we factor $I$ into prime ideals in an ascending order of their norms. We rewrite the factorization as $I=I_{1}I_{2}...I_{s}$, where each of these $I_{j}$’s is a product of the respective prime ideals of the same norm. For example, if $\mathfrak{p}_{i_{1}}^{e_{i_{1}}}\mathfrak{p}_{i_{2}}^{e_{i_{2}}}\supset I_{i}$, for some integers $i,i_{1},i_{2}$, then $N(\mathfrak{p}_{i_{1}})=N(\mathfrak{p}_{i_{2}})$. Therefore,
@@
$=$ $\sum_{j=1}^{s-1}\sum_{\begin{subarray}{c}J\supset I_{j}\\ J\not\supset I_{k^{\prime}},\;\forall k^{\prime}\geq j+1\\ J\neq\mathcal{O}_{K}\end{subarray}}\mu_{K}(J)f(J)\sum_{J^{\prime}\supset I_{j+1}...I_{s}}\mu_{K}(J^{\prime})(\omega_{K}(J)-1+\omega_{K}(J^{\prime})).$ (2.1)
We note that in the above equality, the outer sum on the right ranges to s1s-1, since the contribution from the ideal $I_{s}$ is 0. Also, we do not need the term $f(J^{\prime})$ as by the definition of $f$, we only need to focus on the ideal with the smallest norm, which, of course, contains $J$ and not $J^{\prime}$. Therefore, continuing from (2.1), we have
@@
$+\sum_{j=1}^{s-1}\sum_{\begin{subarray}{c}J\supset I_{j}\\ J\not\supset I_{k^{\prime}},\;\forall k^{\prime}\geq j+1\\ J\neq\mathcal{O}_{K}\end{subarray}}\mu_{K}(J)f(J)\sum_{J^{\prime}\supset I_{j+1}...I_{s}}\mu_{K}(J^{\prime})\omega_{K}(J^{\prime})=:\mathcal{S}_{1}+\mathcal{S}_{2}.$ (2.2)
@@
$(1-x)^{t}=\sum_{i=0}^{t}(-1)^{i}{t\choose i}x^{i}.$ (2.4)
Hence, the RHS of (2.4) with $x=1$ yields the RHS of (2.3), and therefore, we have from both equations that
$\sum_{J^{\prime}\supset I_{j+1}...I_{s}}\mu_{K}(J^{\prime})=\begin{cases}1,\;\text{when $I_{j+1}...I_{s}=\mathcal{O}_{k}$}\\ 0,\;\text{otherwise.}\end{cases}$ (2.5)
From (2.5), it is clear that the only contribution from $\mathcal{S}_{1}$ is when $I_{j+1}...I_{s}=\mathcal{O}_{K}$, which is again only possible if $j=s$. But that is not possible since $j$ ranges from $1$ to s1s-1, and hence, from (2.5), we conclude that
$\mathcal{S}_{1}=0.$ (2.6)
@@
$\sum_{J^{\prime}\supset I_{j+1}...I_{s}}\mu_{K}(J^{\prime})\omega_{K}(J^{\prime})=\sum_{i=1}^{t}(-1)^{i}{t\choose i}i.$ (2.7)
@@
$\mathcal{S}_{2}=-\sum_{\begin{subarray}{c}J\supset I_{s-1}\\ J\not\supset I_{s}\\ J\neq\mathcal{O}_{K}\end{subarray}}\mu_{K}(J)f(J)=Q_{C}^{2}(I).$ (2.10)
To understand the last equality, we observe that $I_{s-1}$ is a product of prime ideals containing $I$ with the second largest norm. Now, the only positive contribution from the terms in the sum in (2.10) is when we choose $J$ to just be the prime ideals, since: